# Study on data analysis for Ives-Stilwell-type experiments

### v11

@@ Abstract
Ives-Stilwell experiment in 1938 was a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. In this paper, we reveal and solve a long-lasting unrecognized fundamental issue that the data analysis for all those experiments was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. For example, in the Letter [Phys. Rev. Lett. 113, 120405 (2014)] the definition of the measurement accuracy of Doppler effect, given by $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, is not physical because Einstein’s Doppler formula cannot be confirmed even when $\varepsilon=0$ holds. Based on existing data, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect. The proposed data analysis in the paper is quite straightforward, but it is the result of the correct way of thinking based on first principles: (i) Einstein’s Doppler effect refers to the same photon (laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. However, the way of thinking set up in the 1938 Ives-Stilwell experiment was incorrect because it was not consistent with the first principles, although it had been followed by quite a few generations of physicists. Even Einstein seemed unaware of this subtle problem with such a far-reaching impact, which highlights the significance of the way of thinking provided in the present paper for analyzing and confirming Einstein’s Doppler effect.
@@ I Introduction
The relativistic Doppler effect is one of the basic results of Einstein’s special relativity, and it states that a plane light wave or a single laser beam or a single photon in free space, observed in any inertial frames of reference, propagates at the same speed but it may exhibit different frequencies and aberrations r1 . The 1938 Ives-Stilwell experiment r2 was a historic research work for confirming Einstein’s special relativity via the Doppler effect, and various modern types of the experiment r3 ; r4 ; r5 ; r6 ; r7 have been repeated by use of laser technology. The concluding result from an Ives-Stilwell-type experiment is presented in the Letter r7 by an international group of collaborators after 15 years of efforts r8 . However the data analysis method for all those experiments r2 ; r3 ; r4 ; r5 ; r6 ; r7 was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. For example, the definition of the measurement accuracy of Doppler effect in r7 is not physical (see Appendix A). Here, based on existing experimental data r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
@@ III Conclusions and remarks
One might argue that the error reported in the paper simply indicates that $\beta~{}[=0.338,377,22]$ is slightly different from the value $[\beta=0.338]$ given in the Letter r7 . However this is not true. The error is that the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ cannot represent the measurement accuracy of Einstein’s Doppler effect, because Einstein’s Doppler effect refers to the effect of a single laser beam.
@@
The present paper might be another interesting example, with the same experimental data used to form an alternative chain of evidence, leading to a unique and definite conclusion. Although the chain of evidence for single laser beams is quite straightforward, it is the result of the correct way of thinking based on first principles: (i) Einstein’s Doppler effect refers to the same photon (laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. However, the way of thinking set up in the 1938 Ives-Stilwell experiment r2 was incorrect because it was not consistent with the first principles, although it had been followed by quite a few generations of physicists r3 ; r4 ; r5 ; r6 ; r7 . Even Einstein seemed unaware of this subtle problem with such a far-reaching impact, which highlights the significance of the way of thinking provided in the present paper for analyzing and confirming Einstein’s Doppler effect.
@@
Acknowledgment. The author acknowledges the Letter by Botermann et al. r7 without which the present work would not have been possible. Note: Like citing published theoretical formulas (Einstein’s Doppler formula, for example), it is allowable and legitimate to cite the experimental data in r7 to support the theory developed in the present paper, and it is not in violation of any ethics guidelines.
Disclosures. The author declares no conflicts of interest.
Data availability. Data underlying the results presented in this paper are available within the paper and Supplement 1 r13 .

### v14

@@ Abstract
Ives-Stilwell experiment in 1938 is a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. However in this paper, we reveal and solve a fundamental issue that the data analysis for all those experiments is not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect and its associated time dilation have not actually been confirmed. For example, in the Letter [Phys. Rev. Lett. 113, 120405 (2014)] the definition of the measurement accuracy of Doppler effect, given by $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, is not physical because Einstein’s Doppler formula cannot be confirmed even when $\varepsilon=0$ holds (see Appendix A). We argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in different inertial frames, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, the data analysis for the 1938 Ives-Stilwell experiment does not comply with the first principles, so it fails to confirm the relativistic effect (see Appendix B), although this data analysis has been mimicked by quite a few generations of physicists. (It should be emphasized that in fact, Ives-Stilwell data analysis and the experimental test itself both fail to confirm Einstein’s Doppler effect.) Based on the first principles, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect.
@@ I Introduction
The relativistic Doppler effect is one of the basic results of Einstein’s special relativity, and it states that a plane light wave or a single laser beam or a single photon in free space, observed in any inertial frames of reference, propagates at the same speed but it may exhibit different frequencies and aberrations r1 . The 1938 Ives-Stilwell experiment r2 was a historic research work for confirming Einstein’s special relativity via the Doppler effect, and various modern types of the experiment r3 ; r4 ; r5 ; r6 ; r7 have been repeated by use of laser technology. The concluding result from an Ives-Stilwell-type experiment is presented in the Letter r7 by an international group of collaborators after 15 years of efforts r8 . However the data analysis method for all those experiments r2 ; r3 ; r4 ; r5 ; r6 ; r7 was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect and its associated time dilation were not actually confirmed. For example, the definition of the measurement accuracy of Doppler effect in r7 is not physical (see Appendix A).
In this paper, we argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, the way of thinking set up in the 1938 experiment did not comply with the first principles so that even Ives and Stilwell themselves were unaware that their experimental test actually did not support Einstein’s Doppler effect (see Appendix B). Here, based on the above first principles with existing experimental data employed r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
@@ II Accuracies and data analysis
One might argue that according to Eqs. (1) and (10)-(17), the accuracy $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ defined in r7 is derived from the use of combined quantities to cancel or reduce systematic errors — a method commonly employed all over experimental science. However, this is not the case because $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ itself cannot confirm Einstein’s Doppler formula even at an ideal situation of $\varepsilon=0$, which has nothing to do with canceling or reducing systematic errors.
In fact, it is also easy to understand mathematically that $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ cannot confirm Einstein’s Doppler formula. In principle, the validity of Einstein’s formulas $\nu_{1}=\gamma(1-\beta)\nu_{p}$ and $\nu_{2}=\gamma(1+\beta)\nu_{a}$ is a sufficient condition for the holding of $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1=0$, but not a necessary condition. The sufficient and necessary condition for $\varepsilon=0$ is that $\nu_{1}\nu_{2}=\nu_{a}\nu_{p}$ holds. For example, $\nu_{1}=10^{-100}\times\gamma(1-\beta)\nu_{p}$ and $\nu_{2}=10^{100}\times\gamma(1+\beta)\nu_{a}$ also make $\nu_{1}\nu_{2}=\nu_{a}\nu_{p}$ or $\varepsilon=0$ holding, although the $\beta$-dependence on $\nu_{1}$ and $\nu_{p}$ or on $\nu_{2}$ and $\nu_{a}$ is already far from Einstein’s Doppler formula in such a case. From this we can see that $\varepsilon=0$ does not necessarily result from Einstein’s Doppler formula. Therefore, the accuracy $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ defined in r7 cannot confirm Einstein’s Doppler formula itself even at $\varepsilon=0$; see Appendix A for more details.
One might argue that although the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ cannot confirm Einstein’s Doppler formula, it can be used to test Einstein’s time dilation when $\varepsilon=0$ holds. But this is not true, because $\varepsilon$ itself is completely independent of the time dilation factor $\gamma$, and one cannot be sure that $\varepsilon=0$ comes from Einstein’s relativity; for example, $\nu_{1}=10^{-100}\times\gamma(1-\beta)\nu_{p}$ and $\nu_{2}=10^{100}\times\gamma(1+\beta)\nu_{a}$ also make $\varepsilon=0$ holding, as mentioned above. Therefore, the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ defined in r7 cannot be used to test Einstein’s time dilation either.
However, if Einstein’s Doppler formula is confirmed, then its associated time dilation will also be verified. That is because, as we all know, Einstein’s Doppler formula is only different by a time dilation factor from the classic Doppler formula.
@@ III Conclusions and remarks
One might argue that the error reported in the paper simply indicates that $\beta~[=0.338,377,22]$ is slightly different from the value $\beta=0.338]$ given in the Letter r7 . However this is not true. The error is that the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ cannot represent the measurement accuracy of Einstein’s Doppler effect, because Einstein’s Doppler effect refers to the effect of a single laser beam.
@@
The present paper might be another interesting example, with the same experimental data used to form an alternative chain of evidence, leading to a unique and definite conclusion. Although the chain of evidence for single laser beams is quite straightforward, it is the result of the correct way of thinking based on first principles: (i) Einstein’s Doppler effect refers to the same photon (laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. However, the way of thinking set up in the 1938 Ives-Stilwell experiment r2 was incorrect because it was not consistent with the first principles (see Appendix B), although it had been mimicked by quite a few generations of physicists r3 ; r4 ; r5 ; r6 ; r7 . Even Einstein seemed unaware of this subtle problem with such a far-reaching impact, which highlights the significance of the first principles proposed in the present paper for analyzing and confirming Einstein’s Doppler effect.
@@ Appendix A Einstein’s Doppler formula not confirmed at $\varepsilon=0$
Appendix B Ives-Stilwell experimental test failing to confirm Einstein’s Doppler effect
First principles indicates that Einstein’s Doppler effect refers to the same photon exhibiting different frequencies observed in different inertial frames, and only the measurement accuracy of Doppler effect from a single light beam can represent the accuracy of Einstein’s Doppler effect, because only this accuracy is a measure to confirm Einstein’s Doppler formula itself. In this appendix, we will demonstrate that Ives and Stilwell did not provide a data analysis for single light beams in their paper r2 , and so violated the first principles. As a result, they were even unware that the experimental data they provided actually did not support Einstein’s Doppler effect.
[…]
In the following, we first examine the data analysis provided by the authors of Ref. r2 , then we propose a unique data analysis for single light beams tailored particularly for the Ives-Stilwell experimental test, and finally, conclusions are given.
Table 1: Experimental data used in Ives-Stilwell analysis, taken from Table I and Table III of Ref. r2 . $\gamma\beta\lambda_{0}\approx\beta\lambda_{0}$ and $(\gamma-1)\lambda_{0}\approx 0.5\beta^{2}\lambda_{0}$ are, respectively, the true values of the observed quantities $(\Delta\lambda)_{\beta}$ and $\Delta\lambda$. The ion’s velocity is computed from $\beta=(\beta\lambda_{0})/\lambda_{0}$, with $\lambda_{0}=4861$Å the ion’s proper wavelength.
Case/Plate/Line  $\beta$   $\beta\lambda_{0}$(Å)   $(\Delta\lambda)_{\beta}$(Å)   $0.5\beta^{2}\lambda_{0}$(Å)   $\Delta\lambda$(Å)
1/169/H3   $0.002,184,736$ $10.62$ $10.35$ $0.0116$ $0.011$
2/160/H2   $0.002,888,295$ $14.04$ $14.02$ $0.0203$ $0.0185$
3/163/H2   $0.003,147,501$ $15.30$ $15.40$ $0.0238$ $0.0225$
4/170/H2   $0.003,361,448$ $16.34$ $16.49$ $0.0275$ $0.027$
5/165/H3   $0.002,855,380$ $13.88$ $14.07$ $0.0198$ $0.0205$
6/172/H2   $0.003,805,801$ $18.50$ $18.67$ $0.0352$ $0.0345$
7/172/H3   $0.003,096,071$ $15.05$ $15.14$ $0.0233$ $0.0215$
8/177/H2   $0.004,433,244$ $21.55$ $21.37$ $0.0478$ $0.047$
Ives-Stilwell data analysis. The second-order combined Doppler effect is described by
$\frac{1}{2}(\lambda_{b}+\lambda_{f})-\lambda_{0}=(\gamma-1)\lambda_{0},$ (21)
where $\gamma$ is the ion’s Lorentz factor, and $\lambda_{0}$ is its proper wavelength; $\lambda_{b}$ and $\lambda_{f}$ are, respectively, the wavelengths of light emitted by the ions in backward and forward directions, observed in the laboratory frame. If both the dependence of $\lambda_{b}$ and the dependence of $\lambda_{f}$ on $\lambda_{0}$ respectively fulfill Einstein’s Doppler formula, then Eq. (21) holds exactly.
In Ives-Stilwell data analysis r2 , the left side of Eq. (21) given by
$\Delta\lambda\equiv\frac{1}{2}(\lambda_{b}+\lambda_{f})-\lambda_{0}$ (22)
is a quantity to be measured (or observed), which means that all $\lambda_{b}$, $\lambda_{f}$, and $\lambda_{0}$ are measured and they may not exactly satisfy Einstein’s Doppler formula, thus causing a deviation from Eq. (21). To examine the deviation, the right side of Eq. (21) given by
$(\gamma-1)\lambda_{0}\approx\frac{1}{2}\beta^{2}\lambda_{0}\quad(\sim\beta^{2},~\textrm{second-order})$ (23)
is taken as the true value of $\Delta\lambda$, where $\beta$ is the ion’s velocity normalized to vacuum light speed. Thus the quantity or measurement accuracy
$\varepsilon=\frac{\Delta\lambda}{(\gamma-1)\lambda_{0}}-1$ (24)
is a measure to confirm the second-order combined Doppler effect described by Eq. (21), because Eq. (21) exactly holds at $\varepsilon=0$, while $\varepsilon\neq 0$ means a deviation from the second-order combined Doppler effect (instead of a single-beam Doppler effect).
Table 2: Accuracies of first-order and second-order combined Doppler effect, computed from the data in Table 1. The accuracies of first-order effect have a range of $2.6\%<\varepsilon_{\beta}<1.4\%$ while the accuracies of second-order effect have a range of $8.9\%<\varepsilon<3.6\%$. Note that if a true value is larger than the value of its observed quantity, then the accuracy is negative; otherwise it is positive.
Case/Plate/Line  $\varepsilon_{\beta}$    $\varepsilon$
1/169/H3   0.0254-0.0254   0.0517-0.0517
2/160/H2   0.0014-0.0014   0.0887-0.0887
3/163/H2   +0.0065+0.0065   0.0546-0.0546
4/170/H2   +0.0092+0.0092   0.0182-0.0182
5/165/H3   +0.0137+0.0137   +0.0354+0.0354
6/172/H2   +0.0092+0.0092   0.0199-0.0199
7/172/H3   +0.0060+0.0060   0.0773-0.0773
8/177/H2   0.0084-0.0084   0.0167-0.0167
The first-order combined Doppler effect is described by
$\frac{1}{2\cos{\theta}}(\lambda_{b}-\lambda_{f})=\gamma\beta\lambda_{0},\vskip 6.0pt$ (25)
where $\theta=7^{\circ}$ is the observation angle. If the dependence of $\lambda_{b}$ and $\lambda_{f}$ on $\lambda_{0}$ respectively fulfill Einstein’s Doppler formula, then Eq. (25) holds exactly.
In their analysis r2 , the left side of Eq. (25) given by
$(\Delta\lambda)_{\beta}=\frac{1}{2\cos{\theta}}(\lambda_{b}-\lambda_{f})$ (26)
is a quantity to be measured (or observed), which means that measured $\lambda_{b}$, $\lambda_{f}$, and $\lambda_{0}$ may result in a deviation from Eq. (25). To examine the deviation, the right side of Eq. (25) given by
$\gamma\beta\lambda_{0}$ $\approx\Big(1+\frac{1}{2}\beta^{2}\Big)\beta\lambda_{0}$
$\approx\beta\lambda_{0}\quad(\sim\beta,~\textrm{first-order})$ (27)
is taken as the true value of $(\Delta\lambda)_{\beta}$. Thus the quantity or measurement accuracy
$\varepsilon_{\beta}=\frac{(\Delta\lambda)_{\beta}}{\gamma\beta\lambda_{0}}-1$ (28)
is a measure to confirm the first-order combined Doppler effect described by Eq. (25), because Eq. (25) exactly holds at $\varepsilon_{\beta}=0$, while $\varepsilon_{\beta}\neq 0$ means a deviation from the first-order combined Doppler effect (instead of a single-beam Doppler effect).
It should be emphasized that physically, $\varepsilon_{\beta}=0$ or Eq. (25) cannot be used to get the ion’s velocity $\beta$, because Eq. (25), just like Eq. (21), is to be confirmed; otherwise, circular reasoning would result. Nevertheless, Ives-Stilwell analysis r2 seemed to have provided this treatment in column 5 of Table III for $0.5\beta^{2}\lambda_{0}$ as a comparison, where $\beta$ was obtained from $\varepsilon_{\beta}=0\Rightarrow(\Delta\lambda)_{\beta}=\gamma\beta\lambda_{0}\approx\beta\lambda_{0}\Rightarrow\beta\approx(\Delta\lambda)_{\beta}/\lambda_{0}$. But such a treatment is not justified if the purpose is to avoid measuring $\beta$, because it is really a logic fallacy.
The authors of Ref. r2 did not provide measured wavelengths explicitly but they can be obtained from Eqs. (22) and (26), given by
$\lambda_{b}=\lambda_{0}+\Delta\lambda+(\Delta\lambda)_{\beta}\cos{\theta}\quad\textrm{(backward light)},$ (29)
$\lambda_{f}=\lambda_{0}+\Delta\lambda-(\Delta\lambda)_{\beta}\cos{\theta}\quad\textrm{(forward light)}.$ (30)
The above measured wavelengths will be used to conduct the data analysis for single light beams.
The experimental data that Ref. r2 provided are shown in Table 1 and the accuracies derived from the data are shown in Table 2. (Note that the ion’s velocity $\beta$ is derived from $\beta=(\beta\lambda_{0})/\lambda_{0}$, with $(\beta\lambda_{0})$ taken from Table I of Ref. r2 , but $(\beta\lambda_{0})$ there is only rounded to the second decimal place, which may cause numerical errors for $0.5\beta^{2}\lambda_{0}$ given in column 4 in Table III of Ref. r2 . For example, $\beta=(\beta\lambda_{0})/\lambda_{0}=15.30/4861\approx 0.003147501$ for case 3, and we have $0.5\beta^{2}\lambda_{0}\approx 0.0241$, compared to $0.5\beta^{2}\lambda_{0}=0.0238$ given in column 4 in Table III of Ref. r2 . But this error would have no essential effect.)
From Tables 1 and 2, we can see that Ives-Stilwell analysis r2 aimed to check how the observed quantities $\Delta\lambda$ and $(\Delta\lambda)_{\beta}$ are close to their true values $(\gamma-1)\lambda_{0}\approx 0.5\beta^{2}\lambda_{0}$ and $\gamma\beta\lambda_{0}\approx\beta\lambda_{0}$, respectively. According to $\Delta\lambda=0.5\beta^{2}\lambda_{0}$ and $\Delta\lambda=[(\Delta\lambda)_{\beta}]^{2}/(2\lambda_{0})$, as shown in Fig. 13 and Fig. 14 of Ref. r2 , the authors concluded that the results exhibiting in the figures “appear to be a satisfactory confirmation of the Larmor-Lorentz theory [Einstein’s special theory]”. However $\Delta\lambda=0.5\beta^{2}\lambda_{0}$ and $\Delta\lambda=[(\Delta\lambda)_{\beta}]^{2}/(2\lambda_{0})$ were obtained by approximate treatment of the true values: $(\gamma-1)\lambda_{0}\approx 0.5\beta^{2}\lambda_{0}$ and $\gamma\beta\lambda_{0}\approx\beta\lambda_{0}$, as shown in Eqs. (23) and (27), and whether they can be used as a criterion is questionable, which can be clearly seen from the data analysis for single light beams shown below.
Data analysis for single light beams. According to Einstein r1 , the Doppler shifted wavelengths for the two light beams are, respectively, given by
$\lambda_{b-\textrm{sh}}=\gamma(1+\beta\cos{\theta})\lambda_{0}$ $\textrm{for backward light},$ (31)
$\lambda_{f-\textrm{sh}}=\gamma(1-\beta\cos{\theta})\lambda_{0}$ $\textrm{for forward light},$ (32)
where $\theta=7^{\circ}$ is the observation angle; $\lambda_{0}$ is the photon’s wavelength observed in the ion’s rest frame, as mentioned before, while $\lambda_{b-\textrm{sh}}$ or $\lambda_{f-\textrm{sh}}$ is the wavelength of the same photon observed in the laboratory frame, which constitutes one of the first principles: Einstein’s Doppler effect refers to the same photon exhibiting different frequencies observed in different inertial frames.
From above, the measurement accuracies of Einstein’s Doppler effect for single light beams are given by
$\varepsilon_{b}=\frac{\lambda_{b}-\lambda_{b-\textrm{sh}}}{\lambda_{b-\textrm{sh}}}$ $\textrm{for backward light},$ (33)
$\varepsilon_{f}=\frac{\lambda_{f}-\lambda_{f-\textrm{sh}}}{\lambda_{f-\textrm{sh}}}$ $\textrm{for forward light},$ (34)
where $\lambda_{b-\textrm{sh}}$ is the theoretically expected value, taken as the true value of the measured wavelength $\lambda_{b}$, while $\lambda_{f-\textrm{sh}}$ is the true value of the measured $\lambda_{f}$.
The accuracies $\varepsilon_{b}$ and $\varepsilon_{f}$ have clear physical meaning. For example, if $\varepsilon_{b}=0$ holds, then we have $\lambda_{b}=\lambda_{b-\textrm{sh}}$ and the Doppler effect of a single photon described by Einstein’s formula Eq. (31) is exactly confirmed. In general, $\varepsilon_{b}~(\neq 0)$ denotes a deviation from the formula for this photon. Thus the quantity $\varepsilon_{b}$ is a measure to confirm Einstein’s formula, and only the Doppler effect from a single photon can represent Einstein’s Doppler effect, which constitutes the other one of the first principles: the quantity (or accuracy) used as a measure to confirm Einstein’s Doppler effect must be able to confirm Einstein’s Doppler formula itself.
From Eqs.(33) and (34) we have
$\lambda_{b}=(1+\varepsilon_{b})\gamma(1+\beta\cos{\theta})\lambda_{0}$ $\textrm{(backward)},$ (35)
$\lambda_{f}=(1+\varepsilon_{f})\gamma(1-\beta\cos{\theta})\lambda_{0}$ $\textrm{(forward)}.$ (36)
According to above Eqs. (35) and (36), we can draw a useful criterion, stating that a necessary condition for observing the relativistic Doppler effect from the backward or forward beam is that
$\varepsilon_{b}>\varepsilon_{c}\quad\textrm{or}\quad\varepsilon_{f}>\varepsilon_{c}\quad\textrm{with}\quad\varepsilon_{c}=\frac{1}{\gamma}-1$ (37)
must hold, where $\varepsilon_{c}$ is the critical accuracy. That is because, for example, if $(1+\varepsilon_{b})\gamma=1$ holds, then we have $\lambda_{b}=(1+\beta\cos{\theta})\lambda_{0}$ which is the classical Doppler wavelength. But the relativistic wavelength must be larger than its corresponding expected classic wavelength. Thus $(1+\varepsilon_{b})\gamma>1$ must hold for Einstein’s Doppler effect, leading to $\varepsilon_{b}>\varepsilon_{c}$. The same argument applies to $\varepsilon_{f}$. Thus the holding of $\varepsilon_{b}>\varepsilon_{c}$ ($\varepsilon_{f}>\varepsilon_{c}$) is a necessary condition to support Einstein’s Doppler effect for the backward (forward) light beam. This simple criterion will be used to examine Ives-Stilwell experimental test r2 , as shown below.
Inserting Eq. (35) into Eq. (29) and Eq. (36) into Eq. (30), we obtain
$\varepsilon_{b}$ $=\frac{D_{b}}{\gamma(1+\beta\cos{\theta})\lambda_{0}}+\varepsilon_{c},$ (38)
$\varepsilon_{f}$ $=\frac{D_{f}}{\gamma(1-\beta\cos{\theta})\lambda_{0}}+\varepsilon_{c},$ (39)
where $D_{b}$ and $D_{f}$ are the discriminants, given by
$D_{b}=\Delta\lambda+\big[(\Delta\lambda)_{\beta}-\beta\lambda_{0}\big]\cos{\theta},$ (40)
$D_{f}=\Delta\lambda-\big[(\Delta\lambda)_{\beta}-\beta\lambda_{0}\big]\cos{\theta}.$ (41)
Note that (i) in above Eqs. (38) and (39), $D_{b}$, $D_{f}$ and $\lambda_{0}$ can be treated as being dimensionless for simplicity and clarity, because they have the same dimension and show up in the form of $D_{b}/\lambda_{0}$ and $D_{f}/\lambda_{0}$, so the dimension can be considered to be cancelled each other; (ii) no approximate treatment is invoked in obtaining $D_{b}$ and $D_{f}$.
[…]
Inserting the data in Table 1 into Eqs. (40) and (41), we obtain all the values of the discriminants $D_{b}$ and $D_{f}$, as shown in Table 3.
From Table 3, we find that for each of all 8 cases in Ives-Stilwell experimental test, we have either $D_{b}<0$ or $D_{f}<0$ holding, so at least one of the two light beams does not have the relativistic effect predicted by Einstein’s special theory, a result contrary to that claimed by the authors of Ref. r2 , which means that the relation expressions $\Delta\lambda=0.5\beta^{2}\lambda_{0}$ and $\Delta\lambda=[(\Delta\lambda)_{\beta}]^{2}/(2\lambda_{0})$ used as a criterion in r2 is not correct. Thus we can conclude that Ives-Stilwell data analysis failed to confirm Einstein’s Doppler effect.
Table 3: Numerical results of discriminants $D_{b}$ and $D_{f}$, computed from the data in Table 1. If $D_{b}$ ($D_{f}$)<0<0, then $\varepsilon_{b}$ ($\varepsilon_{f}$) $\varepsilon_{c}$, and the measured wavelength $\lambda_{b}$ ($\lambda_{f}$) is less than its classic Doppler wavelength, with no expected relativistic effect occurring.
Case/Plate/Line  $D_{b}$    $D_{f}$
1/169/H3   0.2570-0.2570   +0.2790+0.2790
2/160/H2   0.0014-0.0014   +0.0384+0.0384
3/163/H2   +0.1218+0.1218   0.0768-0.0768
4/170/H2   +0.1759+0.1759   0.1219-0.1219
5/165/H3   +0.2091+0.2091   0.1681-0.1681
6/172/H2   +0.2032+0.2032   0.1342-0.1342
7/172/H3   +0.1108+0.1108   0.0678-0.0678
8/177/H2   0.1317-0.1317   +0.2257+0.2257
Einstein’s relativistic Doppler effect for a plane light wave is described by r1
$\lambda_{\textrm{lab}}=\gamma(1-\hat{\mathbf{n}}\cdot\boldsymbol{\beta})\lambda_{0},$ (42)
where $\lambda_{0}$ is the wavelength observed in the moving frame, such as in the ion-rest frame in the Ives-Stilwell case, $\lambda_{\textrm{lab}}$ is the wavelength observed in the laboratory frame, $\boldsymbol{\beta}$ is the velocity at which the frame moves relatively to the laboratory frame, and $\hat{\mathbf{n}}$ is the unit wave vector. From this we can see that the relativistic Doppler effect exists observed in any directions in the laboratory fame as long as $\boldsymbol{\beta}\neq 0$ or $\gamma=(1-\boldsymbol{\beta}^{2})^{-1/2}>1$ holds. However in the Ives-Stilwell experiment test, out of all 8 cases, each has a light beam that does not have the relativistic effect. For example, for case 8 with $D_{b}=-0.1317<0$ leading to $\varepsilon_{b}=-3.7\times 10^{-5}<\varepsilon_{c}=-9.8\times 10^{-6}$, the backward beam has a measured wavelength of 4882.264882.26Å less than its classic Doppler wavelength of 4882.394882.39Å, without any relativistic effect.
From above, we can conclude that the experimental test itself failed to confirm the existence of Einstein’s Doppler effect.
Conclusions. We have demonstrated that Ives and Stilwell did not provide a data analysis for individual light beams in their paper r2 . Their analysis was based on the combined Doppler effect from two light beams generated by hydrogen ions in canal rays, and took the relation expressions $\Delta\lambda=0.5\beta^{2}\lambda_{0}$ and $\Delta\lambda=[(\Delta\lambda)_{\beta}]^{2}/(2\lambda_{0})$ as a criterion. However this was incorrect, as shown in Table 3. Thus Ives-Stilwell data analysis failed to confirm Einstein’s Doppler effect.
According to first principles, we have proposed a justified unique data analysis to confirm Einstein’s Doppler effect for Ives-Stilwell experimental test r2 . We have strictly derived a simple criterion and two discriminants, given by Eq. (37), and Eqs. (40) and (41), which are tailored to analyze the data provided by the experimental test r2 , finding that out of all 8 cases, there is no one to support the relativistic Doppler effect, as shown in Table 3.
Therefore, we conclude that Ives-Stilwell data analysis and the experimental test itself both failed to confirm Einstein’s Doppler effect.
Acknowledgments
The author acknowledges the paper by Ives and Stilwell r2 and the Letter by Botermann et al. r7 , without which the present work would not have been possible. Note: Like citing published theoretical formulas (Einstein’s Doppler formula, for example), it is allowable and legitimate to cite the experimental data in r2 and r7 to support the theory developed in the present paper, and it is not in violation of any ethics guidelines.
Disclosures
The author declares no conflicts of interest.
Data availability
Data underlying the results presented in this paper are available within the paper and Supplement 1 r13 .